OpenStudy (anonymous):

Divide. Write your answer in simplest form. 9/7 ÷ 7/5 9/7 ÷ 5/7 45/49 correct on my mistakes and help me get the right answer please

OpenStudy (chosenmatt):

hi!

OpenStudy (chosenmatt):

one sec :)

OpenStudy (k_lynn):

\[\frac{ 9 }{ 7 }\div \frac{ 7 }{ 5 }=\frac{ 9 }{ 7 }\times \frac{ 5 }{ 7 }\] So yes, You're right!

OpenStudy (anonymous):

I think it is right but that is just me.

OpenStudy (anonymous):

45/49

OpenStudy (anonymous):

I just made my openstudy account

OpenStudy (anonymous):

yay thank you everyone

OpenStudy (anonymous):

Thanks klin

OpenStudy (anonymous):

and welcome @ICanAwnserQuestions :)

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

woah how did i go from 3 to 7?

OpenStudy (anonymous):

cause you helped her

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

lmao